OpenStudy (anonymous):

To solve the equation m+5=14, you should

OpenStudy (anonymous):

m = 14 -5

OpenStudy (anonymous):

subtract 5 from both sides,subtract from both sides,add 5 to both sides, or add to both sides?

OpenStudy (anonymous):

substract 5 from both sides as in m +5 -5 = 14 - 5 m = 14 -5 and then you do the final operation

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no worries :)